# Concurrent-Learning Based Relative Localization in Shape Formation of Robot Swarms

### v1

@@ VI-B Effectiveness of the Shape Formation Scheme
In the swarm shape formation scenario, a swarm of 50 robots are assigned to form a dart shape. Parameters $r_{\mathrm{sense}}$ and $r_{\mathrm{neigh}}$ are chosen as $r_{\mathrm{sense}}=4$m and $r_{\mathrm{neigh}}=2.5$m. Furthermore, collision avoidance radius is set as $r_{\mathrm{avoid}}=1.8$m. Parameters adopted in the relative localization are $\Delta t=0.01$s, $h=20$, $\lambda_{0}=0.1$, $c_{1}=0.1$, $\alpha=0.5$, $\delta_{t}=1$s. Parameters used in the behavior control are $n_{\mathrm{cell}}=2131$, $r_{i}=0.3$m, $v_{\mathrm{max}}=1$m/s, $w_{0}=6$rad/s, $w_{r}=4$rad/s, $\delta_{0}=0.01$, $\kappa_{1}=10$, $\kappa_{2}=15$, $\kappa_{3}=25$, $\kappa_{4}=2$. The process of shape formation is shown in Fig. 5A. During the shape forming process, the different states of each robot are represented by three colors. Red represents that the robot is in neighbor relative localization state and making relative motion to collect measurement data. According to Algorithms 1 and 2, one can see that the velocity command of the red robot is $v_{i}=v_{i}^{\mathrm{enh}}+v_{i}^{\mathrm{int}}$ or $v_{i}=v_{i}^{\mathrm{enh}}$. Magenta represents that the robot is in shape localization agreement state. In this state, according to Algorithm 1, the velocity command is $v_{i}=0$. Blue represents that the robot is in distributed shape formation state. In this state, according to Algorithm 2, the velocity command is $v_{i}=v_{i}^{\mathrm{form}}+v_{i}^{\mathrm{exp}}+v_{i}^{\mathrm{int}}$.
At time instant $t=0$s, each robot moves and collects measurement data to estimate the relative position to its neighbor robots. As some robots complete the relative localization of their neighboring robots at $t=1$s, the process of the shape localization agreement begins. Although some robots are still in red, they will turn to magenta when they collect sufficient data. After the convergence of the seed robot relative position estimator $\hat{q}_{i}(t)$ for each robot $i$, the robots turn into distributed shape formation control state under the action of the hop-count algorithm at $t=7.6$s. The norm value of the shape localization agreement errors $\tilde{q}_{i,0}$ is shown in Fig. 5B. One can see that with the proposed shape localization agreement protocol and the hop-count based end-state decision method, the agreement errors are less than 0.5cm. During the process of the shape formation, some robots may encounter new neighbors and re-enter the neighbor relative localization state, as show in the snapshot of the robots’ states at $t=13$s. According to the norm value of velocity $v_{i}$ shown in Fig. 5B, the robot swarms are static at $t=40$s, which means the shape formation process is completed. With the proposed methods, the robot swarms form the desired dart shape at $t=50$s.
@@
Figure 5: Shape formation process for a swarm of 50 robots. A. Snapshots during the process of shape formation. B. Shape localization agreement error and velocity of each robot.
Figure 6: Simulation results to evaluate the adaptability of the proposed method given different desired shapes and swarm scale $n$. From top to bottom are capital letter “R”, mathematical sum symbol shape and methane molecule shape. From left to right are the cases of 25, 50, 100, 200 robots.
@@ VI-C Adaptability to swarm scale variants
In the adaptability test scenario, robot swarms with different scales are assigned to form different shapes including the capital letter “R”, mathematical sum symbol and the methane molecule shape. For the robot swarms with scale 25, 50, 100 and 200, random initial positions are generated. Parameters adopted in this scenario are the same with those in Section VI-B. The finial position and the trajectory of each robot are shown in Fig. 6. One can see that our strategy exhibits stable performance in different shapes and swarm scales.
Besides, to evaluate the performance of the proposed strategy, statistical simulation results are provided in Fig. 7 and the following three metrics are considered. Fist, the coverage rate is the proportion of the desired shape that are occupied by robots. Second, the entering rate evaluates the proportion of the robots that enter the desired shape. Third, the uniformity gives the distribution uniformity of the robot swarms in the desired shape. The mathematical definitions of the three performance metrics can be found in our previous work [8], we omit them for brief.
@@ VII Experiment Results
In this section, both indoor and outdoor experimental results are provided to evaluate the effectiveness of the proposed relative localization method and the distributed shape formation strategy.
@@ VII-B Relative Docking
In this experiment, a landmark robot is randomly placed in the experimental field, as shown in Fig. 8B. To verify the effectiveness of the proposed relative localization method, another robot named as docking robot is tasked to precisely move to the position $p*=[0.5,-0.5]^{T}$ relative to the landmark robot. Denote the position of the landmark robot and docking robot as $p_{0}$ and $p_{d}(t)$. In the experiment, the control input of the docking robot is simply chosen as the following P-controller,
$v_{d}(t)=-\kappa(\hat{p}_{d,0}(t)-p^{*}),$ (32)
where $\hat{p}_{d,0}(t)$ denotes the estimated relative position between the docking robot and landmark robot. Other parameters are $\kappa=0.02$, $v_{\mathrm{max}}=0.1$m/s, $\Delta t=0.1$, $h=20$, $\lambda_{0}=0.2$.
In this experiment, NOKOV motion capture system is used to record the real position of the landmark robot and the docking robot. The trajectory of the docking robot is shown in Fig. 8B. The norm value of the estimation error $\tilde{q}_{d}(t)$ is shown in Fig. 8B. One can see that when the measurement data is collected sufficiently, the estimation error fast converges into a small set near 0. The norm value $\lVert e_{d}(t)\rVert$ of the tracking error is shown in Fig. 8B. The tracking error is around 10 cm. This experiment verifies the effectiveness of the proposed localization algorithm.
@@ VII-C Distributed Shape Formation
In the outdoor experiments, a swarm consisted of 6 turtlebot robots are assigned to form an “arrow” shape and a capital letter “T” in an outdoor environment range of $8\times 13.5$ m2. The specific settings and parameter descriptions of the experiment are as follows. The nominal measurement range of the sensor is 500m, and the nominal ranging accuracy is 10cm. To simulate robots that can only obtain local information, we artificially limit the measurement range as $r_{\mathrm{sense}}=6$m. To ensure stable communication between the two robots during relative positioning, $r_{\mathrm{neigh}}$ is chosen as $r_{\mathrm{neigh}}=4$m. Furthermore, collision avoidance is triggered if the inter-robot distance is less than $r_{\mathrm{avoid}}=2.5$m. Parameters adopted in the relative localization and shape agreement are $\Delta t=0.02$s, $h=100$, $\lambda_{0}=0.2$, $c_{1}=0.1$, $\alpha=0.5$, $\delta_{t}=1$s. Parameters used in the behavior based control are $r_{i}=0.5$m, $v_{max}=0.15$m/s, $w_{0}=0.32$rad/s, $w_{r}=0.05$rad/s, $\delta_{0}=0.02$, $\kappa_{1}=10$, $\kappa_{2}=15$, $\kappa_{3}=25$, $\kappa_{4}=2$.
Figure 7: Performance indicators for statistical simulation. From top to bottom are coverage rate, entering rate and distribution uniformity of the robot swarms during the process of shape formation. From left to right are the case of capital letter “R”, mathematical sum symbol shape and methane molecule shape.
In this experiment, the cell number of the two graphical binary graph shapes are $n_{cell,arrow}=4209$ and $n_{cell,T}=4266$. one of the six robots is randomly selected as the seed robot and placed in the center of the experimental site. As shown in the Fig. 8A, each robot is equipped with a LED spotlight, and its motion trajectory is represented by the blue trajectory obtained from delay photography. The whole process of the shape formation is shown in Fig. 8C, the algorithm begins at $t=0$s, all the robots begin to move. At $t=24$s, robots move and collect measurement data. After the shape localization agreement, robots move to form the desired shape at $t=48$s. During the shape formation process, some robots may also localize the new neighboring robots. At $t=72$s, the shape formation process is almost complete. The final position of each robot at $t=120$s is represented by a purple highlight while the trajectory of all the robots are also shown as the blue curves in Fig. 8C.
In fact, the actual experimental field is rugged and the robot does not have shock absorption function, which may cause additional measurement error for the IO. Under these adverse conditions, the swarm consisted of 6 robots can still form the specified shape well. It proves the effectiveness, reliability, and robustness of the relative localization based swarm shape formation algorithm proposed in this paper.
@@ VIII Conclusion
This paper has addressed the relative localization based shape formation problem for massive robot swarms. The motivation stems from applications such as object transport and building firefighting, where external localization infrastructures are unavailable. First, we proposed a concurrent-learning based relative position estimator, enabling inter-robot relative localization without the necessity of the well-known PE condition. Second, to achieve the shape localization agreement across the entire swarms in the absense of global coordinates, we proposed a finite-time consensus based agreement protocol. Third, we devised a novel behavior-based shape formation control strategy that can better exploits the historical measurement information to enhance the observability of relative localization among robots. This control strategy has a concise structure and can be implemented on robotic systems easily. Extensive simulation results and comparisons with state-of-the-art relative localization method were presented. Additionally, outdoor experiments involving up to six robots were conducted to verify the robustness, effectiveness and adaptability of the proposed methods.
Appendix A Preliminary lemmas
To establish the main results, the following lemmas are useful.
@@ Appendix A Preliminary lemmas
[48] For any vector $x=[x_{1},...,x_{n}]^{T}\in\mathbb{R}^{n}$, the $p$-norm of $x$ is defined as $\lVert x\rVert_{p}=\left(\sum_{i=1}^{n}|x_{i}|^{p}\right)^{\frac{1}{p}}$. For the constants $p$ and $q$, if $p\geq q\geq 1$, the following inequality holds:
@@
Case 4: $p_{i}<0$, $p_{i}<q_{i}$, consider the following two sub-cases:
@@
$-p^{T}\mathrm{sig}(p-q)^{\alpha}$ $\leq-\frac{\alpha}{1+\alpha}\sum_{i=1}^{n}\lvert p_{i}\rvert^{1+\alpha}+\frac{\alpha}{1+\alpha}2^{\frac{{1+\alpha}}{\alpha}}\sum_{i=1}^{n}\lvert q_{i}\rvert^{1+\alpha}$
$\leq-\frac{\alpha}{1+\alpha}\lVert p\rVert_{1+\alpha}^{1+\alpha}+\frac{\alpha}{1+\alpha}2^{\frac{{1+\alpha}}{\alpha}}\lVert q\rVert_{1+\alpha}^{1+\alpha}$
$\leq-\frac{\alpha}{1+\alpha}\lVert p\rVert^{{1+\alpha}}+\frac{\alpha}{1+\alpha}2^{\frac{{1+\alpha}}{\alpha}}n^{\frac{1-\alpha}{2(\alpha+1)}}\lVert q\rVert^{1+\alpha}.$ (36)
The last inequality in (A) is according to Lemma 3. ∎
@@
Then the state $x$ of the closed-loop system converges into the set $x\in\left\{V(x)^{\alpha}\leq\frac{\beta}{(1-\gamma)K}\right\}$ in finite time $t^{*}$, where $t^{*}$ is defined as
@@
For symmetric matrices $M$ in the following form,
$M=\begin{bmatrix}a_{1}&a_{2}\\ a_{2}&a_{3}\end{bmatrix},$
where $a_{1}>0$, $a_{3}>0$ and $a_{2}\in\mathbb{R}$. Then the maximum value of the $\frac{\lambda_{\min}(M)}{\lambda_{\min}(M)}=\frac{\min\{a_{1},a_{3}\}}{\max\{a_{1},a_{3}\}}$, if and only if $a_{2}=0$.
@@
It is easy to calculate that
$\lambda_{\min}(M)=\frac{a_{1}+a_{3}-\sqrt{(a_{1}-a_{3})^{2}+4a_{2}^{2}}}{2},$ (37a)
$\lambda_{\max}(M)=\frac{a_{1}+a_{3}+\sqrt{(a_{1}-a_{3})^{2}+4a_{2}^{2}}}{2}.$ (37b)
According to (37), one can see that $\frac{\lambda_{\min}(M)}{\lambda_{\min}(M)}$ is a monotonically decreasing function about $a_{2}^{2}$, $\frac{\lambda_{\min}(M)}{\lambda_{\min}(M)}$ take the maximum value $\frac{\min\{a_{1},a_{3}\}}{\max\{a_{1},a_{3}\}}$ at $a_{2}=0$. ∎
Appendix B Proof of Theorem 1
@@
Choose a Lyapunov candidate as
$V_{a}(t_{k})=\tilde{p}_{ij,0}^{T}(t_{k})\tilde{p}_{ij,0}(t_{k}).$ (39)
Define $\Delta V_{a}\triangleq V_{a}(t_{k+1})-V_{a}(t_{k})$, substituting (38) in (39). Then it has
$\Delta V_{a}=$ $\tilde{p}_{ij,0}^{T}(t_{k})\left(I-\eta U_{ij}(t_{k})-\eta S_{ij}\right)^{2}\tilde{p}_{ij,0}(t_{k})$ (40)
$-\tilde{p}_{ij,0}^{T}(t_{k})\tilde{p}_{ij,0}(t_{k})$
$=$ $\tilde{p}_{ij,0}^{T}(t_{k})\left(-2\eta S_{ij}+2\eta^{2}S_{ij}U_{ij}(t_{k})+\eta^{2}S_{ij}^{2}\right)\tilde{p}_{ij,0}(t_{k})$
$-\tilde{p}_{ij,0}^{T}(t_{k})\eta U_{ij}(t_{k})(2I-\eta U_{ij}(t_{k}))\tilde{p}_{ij,0}(t_{k}).$
@@
$\lambda_{\min}(U_{ij}(t_{k}))\lVert\tilde{p}_{ij,0}(t_{k})\rVert\leq\lVert U_{ij}\tilde{p}_{ij,0}(t_{k})\rVert,$ (41a)
$\lambda_{\max}(U_{ij}(t_{k}))\lVert\tilde{p}_{ij,0}(t_{k})\rVert\geq\lVert U_{ij}\tilde{p}_{ij,0}(t_{k})\rVert,$ (41b)
$\lambda_{\min}(S_{ij})\lVert\tilde{p}_{ij,0}(t_{k})\rVert\leq\lVert S_{ij}\tilde{p}_{ij,0}(t_{k})\rVert,$ (41c)
$\lambda_{\max}(S_{ij})\lVert\tilde{p}_{ij,0}(t_{k})\rVert\geq\lVert S_{ij}\tilde{p}_{ij,0}(t_{k})\rVert.$ (41d)
Substituting (41) in (40), it has
$\Delta V_{a}\leq$ $\left(-2\eta\lambda_{\min}(S_{ij})+2\eta^{2}\lambda_{\max}(U_{ij}(t_{k}))\lambda_{\max}(S_{ij})\right.$ (42)
$\left.+\eta^{2}\lambda_{\max}(S_{ij})^{2}\right)\lVert\tilde{p}_{ij,0}(t_{k})\rVert^{2}$
$+\eta^{2}\lambda_{\max}(U_{ij}(t_{k}))^{2}\lVert\tilde{p}_{ij,0}(t_{k})\rVert^{2}.$
With the learning rate $\eta$ (8), (42) can be simplified as
$\Delta V_{a}\leq$ $\frac{-\lambda_{\min}(S_{ij})^{2}}{(\lambda_{\max}(U_{ij}(t_{k}))(t_{k})+\lambda_{\max}(S_{ij}))^{2}}\lVert\tilde{p}_{ij,0}(t_{k})\rVert^{2}.$ (43)
One can see that for matrix $U_{ij}(t_{k})=u_{ij}(t_{k})u_{ij}(t_{k})^{T}\in\mathbb{R}^{2}$ in (43), the maximum eigenvalue satisfies $\lambda_{\max}(U_{ij}(t_{k}))=\lVert u_{ij}(t_{k})\rVert\leq 2v_{\max}\Delta t$. Then (43) can be simplified as

### v4

@@ VI-B Effectiveness of the Shape Formation Scheme
Figure 6: Shape formation process for a swarm of 50 robots. A. Snapshots of the shape formation process. From left to right and from top to bottom are the representative moments of the robots’ states, positions, and motion trajectories. B. Shape localization agreement error and velocity command of each robot.
In the swarm shape formation scenario, a swarm of 50 robots is assigned to form a dart shape. Compared with the latest result [7] and our previous work [5], only onboard UWB and IO measurements are used to achieve the shape formation task for robot swarms. The parameters $r_{\mathrm{sense}}$ and $r_{\mathrm{neigh}}$ are chosen as $r_{\mathrm{sense}}=4$m and $r_{\mathrm{neigh}}=2.5$m. Furthermore, the collision avoidance radius is set as $r_{\mathrm{avoid}}=1.8$m. The localization enhancement motion radius and angle velocity are $r_{i}=0.3$m, $v_{\mathrm{max}}=1$m/s, $w_{0}=6$rad/s, $w_{r}=4$rad/s. The parameters adopted in the relative localization are $\Delta t=0.05$s, $h=40$, $\lambda_{0}=0.1$, $c_{1}=0.1$, $\alpha=0.5$, $\delta_{t}=1$s, $\delta_{0}=0.01$m. The parameters used in the behavior control are $n_{\mathrm{cell}}=2130$, $\kappa_{1}=10$, $\kappa_{2}=15$, $\kappa_{3}=25$, $\kappa_{4}=2$. For more information about parameter selection, please refer to Appendix B-B.
The process of shape formation is shown in Fig. 6A. During the shape forming process, the different states of each robot are represented by three colors. Red represents that the robot is in neighbor relative localization state and making relative motion to collect measurement data. According to Algorithms 1 and 2, one can see that the velocity command of the red robot is $v_{i}=v_{i}^{\mathrm{enh}}+v_{i}^{\mathrm{int}}$ or $v_{i}=v_{i}^{\mathrm{enh}}$. Magenta represents that the robot is in shape localization agreement state. In this state, according to Algorithm 1, the velocity command is $v_{i}=0$. Blue represents that the robot is in distributed shape formation state. In this state, according to Algorithm 2, the velocity command is $v_{i}=v_{i}^{\mathrm{form}}+v_{i}^{\mathrm{exp}}+v_{i}^{\mathrm{int}}$.
At time instant $t=0$s, each robot moves and collects measurement data to estimate the relative position to its neighbor robots. As some robots complete the relative localization of their neighboring robots at $t=1$s, the process of the shape localization agreement begins. Although some robots are still in red, they will turn to magenta when they collect sufficient data. After the convergence of the seed robot relative position estimator $\hat{q}_{i}(t)$ for each robot $i$, the robots turn into distributed shape formation control state under the action of the hop-count algorithm at $t=7.6$s. The norm value of the shape localization agreement errors $\tilde{q}_{i,0}$ is shown in Fig. 6B. One can see that with the proposed shape localization agreement protocol and the hop-count based end-state decision method, the agreement errors are less than 0.5cm. During the process of the shape formation, some robots may encounter new neighbors and re-enter the neighbor relative localization state, as shown in the snapshot of the robots’ states at $t=13$s. It can be seen that the desired shape is almost formed by the robot swarms at $t=30$s. According to the norm value of velocity $v_{i}$ shown in Fig. 6B, the robot swarms are static at $t=40$s, which means the shape formation process is completed. With the proposed methods, the robot swarms form the desired dart shape at $t=50$s.
@@ VI-C Adaptability to Swarm Scale Variants
In the adaptability test scenario, robot swarms with different scales are assigned to form the capital letter “R” shape. For the robot swarms with scale 25, 50, 100 and 200, random initial positions are generated. The parameters adopted in this scenario are the same as those in Section VI-B. The finial position and the trajectory of each robot are shown in Fig. 7 A. One can see that our strategy exhibits stable performance in different shapes and swarm scales.
@@ VII Experiment Results
In this section, indoor and outdoor experimental results are provided to evaluate the effectiveness of the proposed relative localization method and distributed shape formation strategy. The video of the simulation and experiment can be found in https://youtu.be/qMuPmzggDeU.
@@ VII-A Experiment Setup
Figure 7: Simulation results to evaluate the adaptability of the proposed method given different swarm scale $n$. A. Trajectory and final position of the swarm with 25, 50, 100 and 200 robots. B. Coverage rate, entering rate and distribution uniformity of the robot swarms to form capital letter “R”.
@@ VII-B Relative Docking
In this experiment, a landmark robot is randomly placed in the experimental field, as shown in Fig. 8B. To verify the effectiveness of the proposed relative localization method, another robot named as docking robot is tasked to precisely move to the position $p*=[0.5,-0.5]^{T}$ relative to the landmark robot. Denote the position of the landmark robot and docking robot as $p_{0}$ and $p_{d}(t)$. In the experiment, the control input of the docking robot is simply chosen as the following P-controller $v_{d}(t)=-\kappa(\hat{p}_{d,0}(t)-p^{*})$, where $\hat{p}_{d,0}(t)$ denotes the estimated relative position between the docking robot and landmark robot. Other parameters are $\kappa=0.02$, $v_{\mathrm{max}}=0.1$m/s, $\Delta t=0.1$, $h=20$, $\lambda_{0}=0.2$.
In this experiment, NOKOV motion capture system is used to record the real position of the landmark robot and the docking robot. The trajectory of the docking robot is shown in Fig. 8B. The norm value of the estimation error $\tilde{q}_{d}(t)$ is shown in Fig. 8B. One can see that when the measurement data is collected sufficiently, the estimation error fast converges into a small set near 0. The norm value $\lVert e_{d}(t)\rVert$ of the tracking error is shown in Fig. 8B. The tracking error is around 10 cm, which verifies the effectiveness of our localization algorithm.
@@ VII-C Distributed Shape Formation
In the outdoor experiments, a swarm consisting of 6 turtlebot robots is assigned to form an “arrow” shape and a capital letter “T” in an outdoor environment range of $8\times 13.5$ m2. The specific settings and parameter descriptions of the experiment are as follows. The nominal measurement range of the sensor is 500m, and the nominal ranging accuracy is 10cm. To simulate robots that can only obtain local information, we artificially limit the measurement range as $r_{\mathrm{sense}}=6$m. To ensure stable communication between the two robots during relative positioning, $r_{\mathrm{neigh}}$ is chosen as $r_{\mathrm{neigh}}=4$m. Furthermore, collision avoidance is triggered if the inter-robot distance is less than $r_{\mathrm{avoid}}=2.5$m. The parameters adopted in the relative localization and shape agreement are $\Delta t=0.05$s, $h=40$, $\lambda_{0}=0.2$, $c_{1}=0.1$, $\alpha=0.5$, $\delta_{t}=1$s, $\delta_{0}=0.02$. The parameters used in the behavior based control are $r_{i}=0.5$m, $v_{max}=0.15$m/s, $w_{0}=0.32$rad/s, $w_{r}=0.05$rad/s, $\kappa_{1}=10$, $\kappa_{2}=15$, $\kappa_{3}=23$, $\kappa_{4}=2$.
In this experiment, the cell number of the two graphical binary graph shapes are $n_{\mathrm{cell,arrow}}=2209$ and $n_{\mathrm{cell,T}}=2266$. One of the six robots is randomly selected as the seed robot and placed in the center of the experimental site. As shown in the Fig. 8A, each robot is equipped with a LED spotlight, and its motion trajectory is represented by the blue trajectory obtained from delay photography. The whole process of the shape formation is shown in Fig. 8C, the algorithm begins at $t=0$s, all the robots begin to move. At $t=24$s, robots move and collect measurement data. After the shape localization agreement, robots move to form the desired shape at $t=48$s. During the shape formation process, some robots may also localize the new neighboring robots. At $t=72$s, the shape formation process is almost complete. The final position of each robot at $t=120$s is represented by a purple highlight while the trajectories of all the robots are also shown as the blue curves in Fig. 8C. In fact, the actual experimental field is rugged and the robot does not have a shock absorption function, which may cause additional measurement errors for the IO. Under these adverse conditions, the swarm consisting of 6 robots can still form the specified shape well. It proves the effectiveness, reliability, and robustness of the relative localization based shape formation algorithm.
@@ VIII Conclusion
This article has addressed the relative localization based shape formation problem for massive robot swarms. The motivation stems from applications such as object transport and building firefighting, where external localization infrastructures are unavailable. First, we proposed a concurrent-learning based relative position estimator, enabling inter-robot relative localization without the necessity of the well-known PE condition. Second, to achieve the shape localization agreement across the entire swarms in the absence of global coordinates, we proposed a finite-time consensus based agreement protocol. Third, we devised a novel behavior-based shape formation control strategy that can better exploit the historical measurement information to enhance the observability of relative localization among robots. This control strategy has a concise structure and can be implemented on robotic systems easily. Extensive simulation results and comparisons with state-of-the-art relative localization methods were presented. Additionally, outdoor experiments involving up to six robots were conducted to verify the robustness, effectiveness and adaptability of the proposed methods.
Appendix A Theoretical Analysis
A-A Preliminary lemmas
The following lemmas are useful to establish main results.
@@ A-A Preliminary lemmas
[44] For any vector $x=[x_{1},...,x_{n}]^{T}\in\mathbb{R}^{n}$, the $p$-norm of $x$ is defined as $\lVert x\rVert_{p}=\left(\sum_{i=1}^{n}|x_{i}|^{p}\right)^{\frac{1}{p}}$. For the constants $p$, $q$, if $p\geq q\geq 1$, the following inequality holds:
@@
Case 4: $p_{i}<0$, $p_{i}<q_{i}$, consider the two sub-cases:
@@
$-p^{T}\mathrm{sig}(p-q)^{\alpha}\leq-\frac{\alpha}{1+\alpha}\lVert p\rVert^{{1+\alpha}}+\frac{\alpha}{1+\alpha}2^{\frac{{1+\alpha}}{\alpha}}n^{\frac{1-\alpha}{2(\alpha+1)}}\lVert q\rVert^{1+\alpha}.$
@@
Then the state $x$ of the closed-loop system converges into the set $x\in\left\{V(x)^{\alpha}\leq\frac{\beta}{(1-\gamma)K}\right\}$ in finite time $t^{*}$, where $t^{*}$ is
@@
For symmetric matrices $M=\begin{bmatrix}a_{1}&a_{3}\\ a_{3}&a_{2}\end{bmatrix}$, where $a_{1}>0$, $a_{2}>0$ and $a_{3}\in\mathbb{R}$. Then the maximum value of the $\frac{\lambda_{\min}(M)}{\lambda_{\min}(M)}=\frac{\min\{a_{1},a_{2}\}}{\max\{a_{1},a_{2}\}}$, if and only if $a_{3}=0$.
@@
One can calculate that $\lambda_{\min}(M),\lambda_{\max}(M)=\frac{a_{1}+a_{2}\pm\sqrt{(a_{1}-a_{2})^{2}+4(a_{3})^{2}}}{3}.$ One can see that $\frac{\lambda_{\min}(M)}{\lambda_{\min}(M)}$ is a monotonically decreasing function about $(a_{3})^{2}$, $\frac{\lambda_{\min}(M)}{\lambda_{\min}(M)}$ take the maximum value $\frac{\min\{a_{1},a_{2}\}}{\max\{a_{1},a_{2}\}}$ at $a_{3}=0$. ∎
A-B Proof of Theorem 1
@@
Choose a Lyapunov candidate as $V_{a}(t_{k})=\tilde{p}_{ij,0}^{T}(t_{k})\tilde{p}_{ij,0}(t_{k}).$ Define $\Delta V_{a}\triangleq V_{a}(t_{k+1})-V_{a}(t_{k})$, then it has,
$\Delta V_{a}=$ $\tilde{p}_{ij,0}^{T}(t_{k})\left(-2\eta S_{ij}+2\eta^{2}S_{ij}U_{ij}(t_{k})+\eta^{2}S_{ij}^{2}\right)\tilde{p}_{ij,0}(t_{k})$
$-\tilde{p}_{ij,0}^{T}(t_{k})\eta U_{ij}(t_{k})(2I-\eta U_{ij}(t_{k}))\tilde{p}_{ij,0}(t_{k}).$ (34)
@@
$\lambda_{\min}(U_{ij}(t_{k}))\lVert\tilde{p}_{ij,0}(t_{k})\rVert\leq\lVert U_{ij}\tilde{p}_{ij,0}(t_{k})\rVert$ (35a)
$\lambda_{\max}(U_{ij}(t_{k}))\lVert\tilde{p}_{ij,0}(t_{k})\rVert\geq\lVert U_{ij}\tilde{p}_{ij,0}(t_{k})\rVert.$ (35b)
Similar to $S_{ij}$, substituting (35) in (A-B), it has
$\Delta V_{a}$ $\leq\lVert\tilde{p}_{ij,0}(t_{k})\rVert^{2}\left(-2\eta\lambda_{\min}(S_{ij})+2\eta^{2}\lambda_{\max}(U_{ij}(t_{k}))\lambda_{\max}(S_{ij})\right.$
$\left.+\eta^{2}\lambda_{\max}(S_{ij})^{2}-2\eta\lambda_{\min}(U_{ij}(t_{k}))+\eta^{2}\lambda_{\max}(U_{ij}(t_{k}))^{2}\right).$
With the learning rate $\eta$ defined in (7), it has
$\Delta V_{a}\leq$ $\frac{-\lambda_{\min}(S_{ij})^{2}}{(\lambda_{\max}(U_{ij}(t_{k}))(t_{k})+\lambda_{\max}(S_{ij}))^{2}}\lVert\tilde{p}_{ij,0}(t_{k})\rVert^{2}.$
One can see that the maximum eigenvalue of $U_{ij}(t_{k})$ satisfies $\lambda_{\max}(U_{ij}(t_{k}))=\lVert u_{ij}(t_{k})\rVert\leq 2v_{\max}\Delta t$. Then it has,